OpenStudy (anonymous):

Find Area: 2 - √5 x 8 - √5 Answers: 21 41 21 - 10√5 41 - 10√5

OpenStudy (jdoe0001):

\(\bf (2-\sqrt{5})(8-\sqrt{5}) \implies 2(8-\sqrt{5})-\sqrt{5}(8-\sqrt{5})\)

OpenStudy (anonymous):

??

OpenStudy (anonymous):

the fourth one?

OpenStudy (jdoe0001):

dunno, multiply them

OpenStudy (jdoe0001):

let's try the first one \(\bf 2(8-\sqrt{5})\) what does that give you?